OpenStudy (anonymous):

A right square pyramid has base perimeter 62.4m and height 6.4m. Calculate the volume of the pyramid to the nearest cubic metre.

jhonyy9 (jhonyy9):

15,6 m lenght is one side so than the area of base is 15,6*15,6 =243,36 m squared so and the volume V=(A*h)/3 = (243,36*6,4)/3 =1557,504 m cubed